# Quantum Generative Diffusion Model: A Fully Quantum-Mechanical Model for Generating Quantum State Ensemble

### v2

@@ IV Resource-Efficient Quantum Generative Diffusion Model (RE-QGDM)
where $\tau_{t}$ is given by Eq. (6) and $\rho_{t}^{{}^{\prime}}$ is calculated by tracing out the first two qubits of the output from the $U_{1}$ circuit:
@@
Therefore, the trainable parameters for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ of RE-QGDM are $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}\}$. Both QGDM and RE-QGDM share the same diffusion process, training, and generation pipeline, but they differ in the denoising process.
V Critical Evaluation of Denoising in Quantum Generative Diffusion Models
In this section, we provide a detailed analysis of QGDM’s denoising process and present numerical simulation results to support our arguments.
A straightforward design for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ involves using a trainable circuit $U(\boldsymbol{\theta})$ that acts on the extended system $\tau_{t}\otimes\rho_{t}$ and then traces out the qubits of $\tau_{t}$, which can be described by the formula:
$\text{tr}_{(\text{register of }\tau_{t})}\left(U(\boldsymbol{\theta})\left(\tau_{t}\otimes\rho_{t}\right)U^{\dagger}(\boldsymbol{\theta})\right)=\tilde{\rho}_{t-1}.$ (13)
@@ V Critical Evaluation of Denoising in Quantum Generative Diffusion Models
The numerical simulation results illustrating the limitations of this design approach are shown in Fig. 7. Part (a) displays the decreasing trend of the training loss over epochs. Part (b) displays the Hilbert-Schmidt (HS) distance [35, 56] between the input and output states of the denoising circuit $U(\boldsymbol{\theta})$ during training. HS distance, a measure used in quantum information to quantify the distance between two quantum states, is defined as $D_{\text{HS}}(\rho,\sigma)=\sqrt{\text{tr}[(\rho-\sigma)^{2}]}$. Part (c) displays the trend of generation fidelity of QGDM over timestep $t$. The orange solid line represents the fidelity of the generated states with the target state, while the blue dashed line represents the theoretical fidelity curve that the model should achieve. This result shows the model fails to change the input state during the generation process.
Fig. 7 (a) and (b) indicate that with increasing epochs, both training loss and HS distance rapidly decrease. A reduction in HS distance shows that the difference between the input state $\tau_{t}\otimes\rho_{t}$ and the output state $U(\boldsymbol{\theta})(\tau_{t}\otimes\rho_{t})U^{\dagger}(\boldsymbol{\theta})$ becomes minimal, ultimately converging to zero. At this point, tracing over the subsystem containing $\tau_{t}$ would yield a result close to $\rho_{t}$. Fig. 7 (c) shows that QGDM cannot alter the input state during the generation process. The denoising process merely outputs the quantum state fed into it, hence the fidelity between the generated quantum state and the target remains unchanged.
Figure 7: (a) Training loss with respect to the iterations of training. (b) Hilbert-Schmidt (HS) distance between the input and output state of denoising circuit $U(\boldsymbol{\theta})$. (c) After training, the generation fidelity of QGDM with respect to the timestep. The blue dashed line represents the theoretical fidelity that QGDM should achieve, showing the change in the generation process fidelity over timestep $t$. The orange solid line represents the actual generation fidelity.
The comparison between the final generated quantum state $\tilde{\rho}_{0}$ (shown in the first row) and the target quantum state $\rho_{0}$ (displayed in the second row) is illustrated in Fig. 8. Given the complex nature of the density matrix, the figure separately draws the real and imaginary parts, with the former on the left and the latter on the right. This visualization clearly shows that $\tilde{\rho}_{0}$ remains equal to the completely mixed state $\rho_{T}$, indicating that $U(\boldsymbol{\theta})$ essentially reproduces the input state. This highlights the critical need for a well-designed denoising process in QGDM that effectively integrates information between the noisy input state $\rho_{t}$ and the timestep embedding state $\tau_{t}$, especially considering the minor differences in quantum states across adjacent timesteps. The results and discussions in our study will provide valuable insights for the future development of similar quantum machine learning models.
@@
Finally, we observe that $U(\boldsymbol{\theta})$ is not necessarily optimized to the identity matrix $\mathbb{I}$ every time, although it seems straightforward to use $\mathbb{I}$ as a simple way to keep input and output states the same. Learning algorithms do not always choose to do so. This is because different ensembles of quantum states can be transformed into each other by $U(\boldsymbol{\theta})$ while maintaining the same density matrix. There may be more than one transform in the hypothesis space that can satisfy this condition. After training, which one of them is chosen depends on the position of $U(\boldsymbol{\theta})$ in the hypothesis space when the parameters $\boldsymbol{\theta}$ are initialized.
VI Numerical Simulations
@@
Our evaluation of the proposed models involved generating a variety of random quantum states, encompassing both pure and mixed states. We conducted a comparative analysis of our models in comparison to QuGAN [13] and EQ-GAN [21]. QuGAN implementation follows the original framework in Pennylane [70], whereas EQ-GAN is built using the official source code [71].
VI-A Setup
@@
We use the Tensorcircuit framework [72] to simulate quantum circuits and the Tensorflow framework [73] for parameter optimization, with the optimizer Adam [74]. During training, a larger learning rate may make it difficult for the model to converge, while setting the learning rate too small makes the model converge too slowly. To better optimize the model, we use a cosine decay method [75] to dynamically adjust the learning rate. The hyperparameters of QGDM and RE-QGDM are shown in Appendix A.
@@
Our numerical simulations covered a number of qubits ranging from 1 to 8. QGDM achieved results up to $N=4$. EQ-GAN achieved $N=7$ in generating mixed states, as simulating the $N=8$ task requires an unattainable amount of computer memory. RE-QGDM degenerates into QGDM at $N=1$. Therefore, in the next two subsections, the results for QGDM at $N>4$, RE-QGDM at $N=1$, and EQ-GAN at $N=8$ in mixed-state generation are excluded. All simulations were repeated 10 times with different random seeds to obtain the statistical characteristics of our results.
VI-A1 Target Pure State Preparation
VI-A2 Target Mixed State Preparation
Quantum mixed states are generated through $\rho=\sum_{i}{p_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|}$, where $\{\left|\psi_{i}\right\rangle\}$ are randomly produced by the circuit shown in Fig. 9. The probability vector $\{p_{i}\}$ is obtained by uniform sampling from $(0,1]$ and normalized through a softmax operation to ensure $\sum_{i}{p_{i}}=1$ and $p_{i}>0$ for any index $i$. In this paper, we generate a target mixed state by probabilistically mixing two random pure states.
VI-B Pure State Generation
VI-B1 The Impact of $N_{\text{time}}$
@@
We first investigated the impact of $N_{\text{time}}$ in QGDM for generating pure states. The results are shown in Fig. 10. Generation fidelity improves with $N_{\text{time}}$ when $N_{\text{time}}\leq N$, achieving a maximum when $N_{\text{time}}=N$ for all $N$.
Figure 10: The generation fidelity of QGDM in generating target pure states for different qubits with respect to $N_{\text{time}}$. All data represent the average of 10 parallel experiments, and the error bars indicate the standard deviation.
However, when $N_{\text{time}}>N$, except for the case $N=1$, fidelity decreases slightly in other cases. The reason for this may be that there is a redundant qubit that complicates the denoising process. The denoising circuit aims to predict the denoising state $\tilde{\rho}_{t}$ by merging information from $\tau_{t}$ and $\rho_{t}$, and then projecting its prediction onto the first $N$ qubits of the $\tau_{t}$ state. These qubits are denoted as register A in Fig. 3. In cases where $N_{\text{time}}>N$, there is an extra qubit in the state $\tau_{t}$ compared to $\rho_{t}$. This additional qubit may be redundant, complicating the process of merging and output prediction.
When $N_{\text{time}}<N$, register A contains a part of the qubits of $\rho_{t}$. However, the presence of overlapping qubits between register A and the qubits of $\rho_{t}$ may diminish the efficacy of $U(\boldsymbol{\theta})$ in such scenarios since a part of the information from $\rho_{t}$ is already stored in register A. In Section V, we investigate a well-intentioned but faulty denoising design. This design involves transferring all qubits from $\rho_{t}$ to the output register, maximizing the overlap between register A and the qubits of $\rho_{t}$. This extreme case strengthens our hypothesis, illustrating the potential implications.
Based on the results in Fig. 10, subsequent simulations default to $N_{\text{time}}=N$ as this setting yields the best generative results with our proposed method.
VI-B2 Pure States Generation Results
@@
TABLE II: Comparison of pure state generation between QuGAN, EQ-GAN, and our proposed methods.
@@
RE-QGDM (Ours) / 0.994$\pm$8e-3 0.999$\pm$6e-4 0.999$\pm$4e-4 0.997$\pm$4e-3 0.995$\pm$1e-2 0.994$\pm$4e-3 0.990$\pm$1e-2
@@
In Fig. 11, we compare our proposed QGDM and RE-QGDM with QuGAN [13] and EQ-GAN [21] to assess their fidelity in generating the desired pure state. The results indicate that as the number of qubits in the target state increases, the fidelity of all models decreases.
The QuGAN model maintains high fidelity at lower qubit counts. However, its median fidelity decreases as the number of qubits increases, along with an increase in variance. For instances with $N\geq 7$, QuGAN’s fidelity falls below 0.99 for all 10 parallel experiments, indicating its limitations in handling more complex quantum states. The EQ-GAN model achieves a perfect fidelity of 1.0 when the system has only one qubit. However, its fidelity decreases as the number of qubits grows, displaying the largest variance among the four models. From 3 qubits onward, EQ-GAN’s fidelity consistently stays below 0.99, highlighting its challenges in generating pure states. The fidelity of our proposed methods, QGDM and RE-QGDM, remains mostly above 0.99 across all qubit numbers, with more concentrated data distributions, indicating better scalability. In most scenarios, RE-QGDM surpasses the other three models in terms of median fidelity and demonstrates greater stability at higher qubit counts compared to QuGAN and EQ-GAN. In summary, QGDM excels in smaller quantum systems, while RE-QGDM offers a more scalable solution across a broader range of qubit configurations.
To better evaluate the performance differences among models, we present the generative fidelity values for each model in Table. II. The average fidelity of QuGAN diminishes as the number of qubits increases, starting from 0.999$\pm$5e-5 at $N=1$ to 0.971$\pm$1e-1 at $N=8$. It is worth noting that its fidelity drops below 0.99 for $N\geq 6$, indicating a decline in performance as the quantum system size increases. The fidelity of QuGAN from $N=1$ to $5$ remains above 0.99, showing a downward trend for larger systems, although these values are consistently lower than those achieved by our proposed method. The EQ-GAN model achieves the best initial result, producing fidelities of 1 at $N=1$. However, as the number of qubits increases, EQ-GAN shows the lowest average fidelity among the four models, along with the highest variance. QGDM exhibits stable fidelity, with an average of 0.999 at $N=1,2$, and $3$, and 0.996$\pm$8e-3 at $N=4$. Importantly, the RE-QGDM model consistently demonstrates high fidelity across all numbers of qubits, even achieving 0.99 at $N=8$. This underscores the superior robustness and accuracy of RE-QGDM in pure state generation, highlighting the effectiveness and scalability of our models.
VI-B3 Timestep Embedding State Visualization
Fig. 12 shows selected examples visualizing a single-qubit timestep embedding state sequence $\{\tau_{t}\}_{t=1}^{T}$ after training, with $T=30$. Different colors are used to denote the states in each timestep $t$. Each example demonstrates that $\tau_{t}$ shifts its position on the Bloch sphere surface coherently as $t$ progresses. The density of the states varies as well. Fig. 12 (a) and (b) show a more dispersed distribution on the Bloch sphere surface. Subfigures (c) and (d) illustrate cases where states are more clustered. Subfigures (e) and (f) display an intermediate situation, with the state sequence moderately and coherently distributed over the surface of the Bloch sphere.
@@ VI-B3 Timestep Embedding State Visualization
Compared to qubit encoding methods [58, 59, 60], our approach to timestep embedding, as inspired by Ref. [63], is more expressive and flexible. The qubit encoding method generates states that are equidistantly arranged in a ring on the Bloch sphere surface. This restricts subsequent learning models to only learning from such fixed patterns. In contrast, our approach allows the timestep embedding states to be adaptable, enhancing their multifunctionality. This could potentially decrease the difficulty of learning for subsequent models and improve the overall model’s performance.
VI-C Mixed State Generation
In Fig. 13, we evaluate the fidelity of four quantum generative models for generating mixed states. The results indicate that the fidelity of QuGAN and EQ-GAN decreases as the number of qubits increases, while QGDM and RE-QGDM still maintain high generation fidelity in large-scale tasks. Specifically, QuGAN exhibits a continuous decrease in fidelity with an increase in $N$, sharply falling from above 0.9 at $N=1$ to approximately 0.3 at $N=8$ for most fidelity points and the median. EQ-GAN also shows a significant decrease in fidelity but maintains a better median and data point distribution compared with QuGAN. These findings suggest that EQ-GAN outperforms QuGAN in generating mixed states, but both are difficult to handle the task of generating mixed states with multiple qubits. In contrast, QGDM and RE-QGDM consistently show a median fidelity above 0.99 across all scenarios. QGDM surpasses both QuGAN and EQ-GAN at $N=1$, with most fidelity points above 0.99. For $N=$2, 3, and 4, RE-QGDM has a slight advantage over QGDM, with RE-QGDM having more concentrated data points. Notably, at $N\geq 5$, RE-QGDM is significantly better than QuGAN and EQ-GAN, maintaining all fidelity scores above 0.9.

### v3

@@ 4 Resource-efficient Quantum Generative Diffusion Model (RQGDM)
where $\tau_{t}$ is given by (6) and $\rho_{t}^{{}^{\prime}}$ is calculated by tracing out the first two qubits of the output from the $U_{1}$ circuit:
@@
Therefore, the trainable parameters for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ of RQGDM are $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}\}$. Both QGDM and RQGDM share the same diffusion process, training, and generation pipeline, but they differ in the denoising process.
5 Critical Evaluation of Denoising in QGDM
In this section, we provide a detailed analysis of QGDM’s denoising process and present numerical simulation results to support our analysis results.
[…]
$\text{tr}_{\tau_{t}\text{-subsystem}}\left(U(\boldsymbol{\theta})\left(\tau_{t}\otimes\rho_{t}\right)U^{\dagger}(\boldsymbol{\theta})\right)=\tilde{\rho}_{t-1}.$ (13)
@@ 5 Critical Evaluation of Denoising in QGDM
The numerical simulation results illustrating the limitations of this design approach are shown in Fig. 5. Fig. 5(a) displays the decreasing trend of the training loss over epochs. Fig. 5(b) displays the Hilbert-Schmidt (HS) distance [38, 57] between the input and output states of the denoising circuit $U(\boldsymbol{\theta})$ during training. As a measure used in quantum information to quantify the distance between two quantum states, Hs distance is defined as $D_{\text{HS}}(\rho,\sigma)=\sqrt{\text{tr}[(\rho-\sigma)^{2}]}$. Fig. 5(c) displays the trend of generation fidelity of QGDM over timestep $t$. The orange solid line represents the fidelity of the generated states with the target state, while the blue dashed line represents the theoretical fidelity curve that the model should achieve. This result shows the model fails to change the input state during the generation process.
Figs. 5(a)-(b) indicate that with increasing epochs, both training loss and HS distance rapidly decrease. A reduction in HS distance shows that the difference between the input state $\tau_{t}\otimes\rho_{t}$ and the output state $U(\boldsymbol{\theta})(\tau_{t}\otimes\rho_{t})U^{\dagger}(\boldsymbol{\theta})$ becomes minimal, ultimately converging to zero. At this point, tracing over the subsystem containing $\tau_{t}$ would yield a result close to $\rho_{t}$. Fig. 5(c) shows that QGDM cannot alter the input state during the generation process. The denoising process merely outputs the quantum state fed into it. Hence the fidelity between the generated quantum state and target remains unchanged.
Figure 5: (a) Training loss with respect to the iterations of training. (b) Hilbert-Schmidt (HS) distance between the input and output state of denoising circuit $U(\boldsymbol{\theta})$. (c) After training, the generation fidelity of QGDM with respect to the timestep.
The comparison between the final generated quantum state $\tilde{\rho}_{0}$ (shown in the first row) and $\rho_{0}$ (displayed in the second row) is illustrated in Fig. 6. Given the complex nature of the density matrix, the figure separately draws the real and imaginary parts, with the former on the left and the latter on the right. This visualization clearly shows that $\tilde{\rho}_{0}$ remains equal to the completely mixed state $\rho_{T}$, indicating that $U(\boldsymbol{\theta})$ essentially reproduces the input state. This highlights the critical need for a well-designed denoising process in QGDM that effectively integrates information between the noisy input state $\rho_{t}$ and the timestep embedding state $\tau_{t}$, especially considering the minor differences in quantum states across adjacent timesteps. The results and discussions in our study intend to provide valuable insights for the future development of similar quantum machine learning models.
@@
Finally, we observe that $U(\boldsymbol{\theta})$ is not necessarily optimized to identity matrix $\mathbb{I}$ every time, although it seems straightforward to use $\mathbb{I}$ as a simple way to keep input and output states the same. Learning algorithms do not always choose to do so. This is because different ensembles of quantum states can be transformed into each other by $U(\boldsymbol{\theta})$ while maintaining the same density matrix. There may be more than one transform in the hypothesis space that can satisfy this condition. After training, which one of them is chosen depends on the position of $U(\boldsymbol{\theta})$ in the hypothesis space when the parameters $\boldsymbol{\theta}$ are initialized.
6 Numerical Simulations
@@
Our evaluation of the proposed models involves generating a variety of random quantum states, encompassing both pure and mixed states. We have conducted a comparative analysis of our models, QuGAN [9], and EQ-GAN [21]. QuGAN implementation follows the original framework in Pennylane [66], whereas EQ-GAN is built by using the source code in [67].
For more details on the numerical simulations, including training hyperparameters, target state preparation, and the training and generation metrics of our models, please refer to the supplementary file.
6.1 Setup
@@
We use the Tensorcircuit framework [68] to simulate quantum circuits and the Tensorflow framework [69] for parameter optimization, with the optimizer Adam [70]. During training, a larger learning rate may make it difficult for the model to converge, while setting the learning rate too small makes the model converge too slowly. To better optimize the model, we use a cosine decay method [71] to dynamically adjust the learning rate.
@@
Our numerical simulations cover a number of qubits ranging from 1 to 8. QGDM requires $2N$ qubits to generate an $N$-qubit target state. Therefore, in our simulations, QGDM achieves results for $N\leq 4$. EQ-GAN can generate mixed states for $N\leq 7$, as simulating the $N=8$ task demands an unattainable amount of computer memory. At $N=1$, RQGDM degenerates into QGDM. Therefore, in the next two subsections, the results for QGDM at $N>4$, RQGDM at $N=1$, and EQ-GAN at $N=8$ in mixed-state generation are excluded. All simulations are repeated 10 times with different random seeds to obtain the statistical characteristics of our results.
6.2 Pure State Generation
6.2.1 Impact of $N_{\tau}$
@@
We first investigate the impact of $N_{\tau}$ in QGDM for generating pure states. The results are shown in Fig. 7. Generation fidelity improves with $N_{\tau}$ when $N_{\tau}\leq N$, achieving the maximum when $N_{\tau}=N$ for all $N$.
@@
TABLE II: Comparison of pure state generation between QuGAN, EQ-GAN, and our proposed models.
@@
RQGDM (Ours) / 0.994$\pm$8e-3 0.999$\pm$6e-4 0.999$\pm$4e-4 0.997$\pm$4e-3 0.995$\pm$1e-2 0.994$\pm$4e-3 0.990$\pm$1e-2
Figure 7: The generation fidelity of QGDM in generating target pure states for different qubits with respect to $N_{\tau}$.
However, when $N_{\tau}>N$, except for the case $N=1$, fidelity decreases slightly in other cases. The reason for this may be that there is a redundant qubit that complicates the denoising process. The denoising circuit aims to predict the denoising state $\tilde{\rho}_{t}$ by merging information from $\tau_{t}$ and $\rho_{t}$, and then projecting its prediction onto the first $N$ qubits of the $\tau_{t}$ state. These qubits are denoted as register A in Fig. 3. In cases where $N_{\tau}>N$, there is an extra qubit in the state $\tau_{t}$ compared to $\rho_{t}$. This additional qubit may be redundant, complicating the process of merging and output prediction.
When $N_{\tau}<N$, register A contains a part of the qubits of $\rho_{t}$. However, the presence of overlapping qubits between register A and the qubits of $\rho_{t}$ may diminish the efficacy of $U(\boldsymbol{\theta})$ in such scenarios since a part of the information from $\rho_{t}$ is already stored in A. In Section 5, we have investigated a well-intentioned but faulty denoising design. This design involves transferring all qubits from $\rho_{t}$ to the output register, maximizing the overlap between A and the qubits of $\rho_{t}$. This extreme case strengthens our hypothesis, illustrating the potential implications.
Based on the results in Fig. 7, subsequent simulations default to $N_{\tau}=N$ as this setting yields the best generative results with our proposed method.
6.2.2 Pure States Generation Results
In Fig. 8, we compare QGDM and RQGDM with QuGAN [9] and EQ-GAN [21] to assess their fidelity in generating the desired pure state. The results indicate that as the number of qubits in the target state increases, the fidelity of all models decreases.
QuGAN maintains high fidelity at lower qubit counts. However, its median fidelity decreases as the number of qubits increases, along with an increase in variance. For instances with $N\geq 7$, QuGAN’s fidelity falls below 0.99 for all 10 experiments, indicating its limitations in handling more complex quantum states. EQ-GAN achieves a perfect fidelity of 1.0 when the system has only one qubit. However, its fidelity decreases as the number of qubits grows, displaying the largest variance among the four models. From 3 qubits onward, EQ-GAN’s fidelity consistently stays below 0.99, highlighting its challenges in generating pure states. The fidelity of QGDM and RQGDM remains mostly above 0.99 across all qubit numbers, with more concentrated data distributions, indicating their better scalability than QGAN and EQ-GAN. In most scenarios, RQGDM surpasses the other three models in terms of median fidelity and demonstrates greater stability at higher qubit counts than its peers. In summary, QGDM excels in smaller quantum systems, while RQGDM offers a more scalable solution across a broader range of qubit configurations.
@@ 6.2.2 Pure States Generation Results
To better evaluate the performance differences among models, we present the generative fidelity values for each model in Table. II. The fidelity of QuGAN diminishes as the number of qubits increases, starting from 0.999$\pm$5e-5 at $N=1$ to 0.971$\pm$1e-1 at $N=8$. It is worth noting that its fidelity drops below 0.99 for $N\geq 6$, indicating a decline in performance as the quantum system size increases. The fidelity of QuGAN from $N=1$ to $5$ remains above 0.99, showing a downward trend for larger systems, although these values are consistently lower than those achieved by our proposed method. EQ-GAN achieves the best initial result, producing fidelities of 1 at $N=1$. However, as the number of qubits increases, EQ-GAN shows the lowest average fidelity among the four models, along with the highest variance. QGDM exhibits stable fidelity, with an average of 0.999 at $N=1,2$, and $3$, and 0.996$\pm$8e-3 at $N=4$. Importantly, EQ-GAN consistently demonstrates high fidelity across all numbers of qubits, even achieving 0.99 at $N=8$. This underscores the superior robustness and accuracy of RQGDM in pure state generation, highlighting the effectiveness and scalability of our proposed models.
6.2.3 Timestep Embedding State Visualization
Fig. 9 shows selected examples visualizing a single-qubit timestep embedding state sequence $\{\tau_{t}\}_{t=1}^{T}$ after training, with $T=30$. Different colors are used to denote the states in each timestep $t$. Each example demonstrates that $\tau_{t}$ shifts its position on the Bloch sphere surface coherently as $t$ progresses. The density of the states varies as well. Figs. 9(a)-(b) show a more dispersed distribution on the Bloch sphere surface. Figs. 9(c)-(d) illustrate cases where states are more clustered. Figs. 9(e)-(f) display an intermediate situation, with the state sequence moderately and coherently distributed over the surface of the Bloch sphere.
@@ 6.2.3 Timestep Embedding State Visualization
Compared to qubit encoding methods [59, 60], our approach to timestep embedding, as inspired by [63], is more expressive and flexible. The qubit encoding method generates states that are equidistantly arranged in a ring on the Bloch sphere surface. This restricts subsequent learning models to only learning from such fixed patterns. In contrast, our approach allows the timestep embedding states to be adaptable, enhancing their multifunctionality. This could potentially decrease the difficulty of learning for subsequent models and improve the overall model’s performance.
6.3 Mixed State Generation
Fig. 10 shows the fidelity of four quantum generative models for generating mixed states. The results indicate that the fidelity of QuGAN and EQ-GAN decreases as the number of qubits increases, while QGDM and RQGDM maintain high generation fidelity in large-scale tasks. Specifically, QuGAN exhibits a continuous decrease in fidelity with an increase in $N$, sharply falling from above 0.9 at $N=1$ to approximately 0.3 at $N=8$ for most fidelity points and the median. EQ-GAN also shows a significant decrease in fidelity but maintains a better median and data point distribution than QuGAN. These findings suggest that EQ-GAN outperforms QuGAN in generating mixed states, but both are difficult to handle the task of generating mixed states with multiple qubits. In contrast, QGDM and RQGDM consistently show a median fidelity above 0.99 across all scenarios. QGDM surpasses both QuGAN and EQ-GAN at $N=1$, with most fidelity points above 0.99. For $N=$2, 3, and 4, RQGDM has a slight advantage over QGDM, with RQGDM having more concentrated data points. Notably, at $N\geq 5$, RQGDM is significantly better than QuGAN and EQ-GAN, maintaining all fidelity scores above 0.9.